### v1

@@ Abstract
A problem of reconstructing words from their subwords involves determining the minimum amount of information needed, such as multisets of scattered subwords of a specific length or the frequency of scattered subwords from a given set, in order to uniquely identify a word. In this paper we show that a cyclic word on a binary alphabet can be reconstructed by its scattered subwords of length $\frac{3}{4}n+4$, and for each $n$ one can find two cyclic words of length $n$ which have the same set of scattered subwords of length $\frac{3}{4}n-\frac{3}{2}$.
@@ 1 Introduction
The problem of reconstruction of words by their subwords is an important topic in combinatorics of words and related fields of mathematics and theoretical computer science. There are many variants of this problem: reconstruction of normal or cyclic words, from the sets of scattered or contiguous subwords, with or without multiplicities, restricting to a particular family of words and so on. It is known since more than 40 years that one can reconstruct a word of length $n$ from the set of all its scattered subwords of length $\lfloor\frac{n-1}{2}\rfloor$, and this bound is optimal [1]. The same bound holds for contiguous subwords with multiplicities, while without multiplicities it is $(n-1)$ [2].
The question of reconstruction words of length $n$ by their scattered subwords of length $k$ with multiplicities turned out to be more complicated. One of the first results on this question states that for words over a binary alphabet and for $k\geqslant[\frac{n}{2}]$ we can uniquely reconstruct the word, and for $k<\log_{2}(n)$ we cannot do it [3]. One of the best upper bound of this problem with subword length $\lfloor\frac{16}{7}\sqrt{n}\rfloor$ follows from a result related to polynomials introduced in [4]. One of the best lower bounds is $k=\Omega(e^{\log^{\frac{1}{2}}(n)})$ [5].
Other results about words reconstruction include reconstruction from the numbers of occurrences of scattered factors of some special form. For example, a word $w\in\{a,b\}^{*}$ can be reconstructed from the numbers of occurrences of at most $\min(|w|_{a},|w|_{b})+1$ scattered factors of the form $a^{i}b$, where $|w|_{a}$ is the number of occurrences of the letter $a$ in $w$ [6]. A word can qlso be reconstructed from the numbers of occurrences as scattered subwords of $O(\ell^{2}\sqrt{n\log(n)})$ words we have chosen, where $\ell$ is the cardinality of alphabet [9]. Besides that, there are some results about reconstruction of words of special types. For example, reconstruction of DNA-words has been studied in [7], and reconstruction of words that are cyclic shifts of some fixed word has been explored in [8].
@@
Problem: Consider a cyclic word $w$ on binary alphabet of length $n$ and the set of its scattered subwords (without multiplicities) of length at most $k$. For which $k$ can we always recover $w$?
@@
For any two different cyclic binary words $u$ and $v$ of length $n$ there exists a word $w$ of length at most $\frac{3}{4}n+4$ such that $w$ is a subword of only one of the words $u$ and $v$.
We will also show that for each $n\geqslant 7$ there are pairs of words for which sets of subwords of length at most $\frac{3}{4}n-\frac{3}{2}$ are equal (see Proposition 3.1). So, we found a lower bound $k\geqslant\frac{3}{4}n-\frac{3}{2}$ and an upper bound $k\leqslant\frac{3}{4}n+4$. In other words, we calculated $k$ modulo some constant which is at most $5$, so the bound is almost sharp.
The paper is organized as follows. In Section 2, we introduce the necessary definitions and notation used throughout the text. In Section 3, we prove some auxiliary propositions that are needed for the proof of the theorem. Besides that, in Section 3 we provide a lower bound for $k$. In Section 4, we prove the main theorem. The section is divided into two subsections: in Subsection 4.1 we introduce the notation and the general structure of the proof, and the Subsection 4.2 contains the proof of the theorem divided into five lemmas.
@@ 2 Definitions and notation
Two words $x$ and $y$ are said to be conjugate if there exist words $u$ and $v$ such that $x=uv$ and $y=vu$. We define a cyclic word as an equivalence class of the conjugacy relation on $\Sigma^{*}$. Thus, if $w\in\Sigma^{*}$, then the cyclic word represented by $w$ is the set $\{vu\in\Sigma^{*}\mid u,v\in\Sigma^{*},uv=w\}$. In the further text, when referring to a cyclic word, we will write a representative of this class, slightly abusing the notion to avoid cluttering the text.
Let $w$ be a cyclic word. A subword of the form $0^{+}$ (resp., $1^{+}$) continued to the left and to the right with 1 (resp., 0) is called a block of 0’s (resp., 1’s). For a cyclic word $w$ with a representative $w_{1}\cdots w_{n}$, a subword of $w$ is a cyclic word with a representative of the form $w_{i_{1}}\cdots w_{i_{k}}$, where $1\leqslant i_{1}<\ldots<i_{k}\leqslant n$.
We say that $w$ is periodic if $w=(1^{\alpha_{1}}0^{\beta_{1}}1^{\alpha_{2}}0^{\beta_{2}}\cdots 1^{\alpha_{s}}0^{\beta_{s}})^{r}$ for some $\alpha_{i}\geqslant 1$, $\beta_{i}\geqslant 1$ and $r\geqslant 2$.
Let $w$ be a cyclic word. Throughout the paper, we make use of the following notation:
  • $n_{w}$ is the length of $w$
  • $n_{0,w}$ and $n_{1,w}$ are the numbers of 0’s and 1’s in $w$
  • $2l_{w}$ is the total number of blocks in $w$, $l_{w}$ blocks of 0’s and $l_{w}$ blocks of 1’s.
  • $x_{w}$ is the length of the longest block of 0’s in $w$
@@
  • $y_{w}$ is the length of the longest block in $w$ which is shorter than $0^{x_{w}}$ (we might have $y_{w}=0$).
  • $n_{long,w}$ is the number of 00’s in big blocks and $n_{short,w}=n_{0}-n_{long,w}$ is the number of 00’s in small blocks.
  • $w_{long}$ is the subword of $w$ which contains all 1’s of $w$ and all 0’s from big blocks of 0’s (all blocks $0^{x_{w}}$).
  • $w_{short}$ is the subword of $w$ which contains all 1’s of $w$ and all 0’s from small blocks of 0’s (all blocks of 0’s except for the blocks $0^{x_{w}}$).
  • $w_{zeros}$ is the subword of $w$ which contains all 1’s of $w$ and one 00 from each block of 0’s.
In the further text, we will omit the subscript $w$ when no confusion arises.
@@
Let $w=10^{3}101^{2}0^{3}1^{2}0^{2}$. Then $n=15,n_{1}=6,n_{0}=9,x=3,a=2,n_{long}=6,n_{short}=3,w_{long}=1^{3}0^{3}1^{3}0^{3},w_{zeros}=10101^{2}01^{2}0$.
For the proof of the main result, we will need to treat words of certain specific forms separately. We hence introduce the following definition:
@@
Let $w$ be a word with $l_{w}\geqslant 2$. We say that $w$ is special if it is has one of the following three types:
@@
    $w=0^{t}1^{m}0^{t}1^{m}\cdots 0^{t}1^{m}0^{t}1^{2m}0^{t}1^{m}0^{t}1^{m}\cdots 0^{t}1^{m}0^{t}1^{2m}$
    for some positive integers $t$ and $m$.
In other words, a word of the first type has $l_{w}$ blocks $1^{m}$, a word of the second type has $l_{w}-1$ blocks $1^{m}$ and one block $1^{2m}$, a word of the third type has $l_{w}-2$ blocks $1^{m}$ and two blocks $1^{2m}$, and the number of blocks of 00’s is at least two in each case.
For a special word $w$, we let $dist(w)$ denote the number $m$ from the definition of a special word. In fact, $m$ is the length of the shortest block of 1’s.
When working with cyclic words, sometimes we will need to index their letters, only one letter or both letters. For example, for a cyclic word generated by $00101$ we could either index 1’s as $001_{1}01_{2}$, or as $001_{2}01_{1}$. We remark that it is not exactly the same as choosing a representative from the conjugacy class (the difference comes up in periodic cyclic words).
@@
Let $u$ and $v$ be cyclic words with indexed $1$’s such that $n_{0,u}=n_{0,v}$ and $n_{1,u}=n_{1,v}$. We then define overlay by 1’s of $u$ and $v$ as a bijection between indexed 1’s in $u$ and in $v$ which, for some integer $i$, translates each $1_{j}$ in $u$ to $1_{j+i}$ in $v$ (indices are taken modulo $n_{1}$).
In other words, such overlay can be seen as an order-preserving bijection between 1’s in words $u$ and $v$, or simply as a shift of indices of 1’s in $v$ relatively $u$. This bijection between 1’s induces a bijection between blocks of 00’s in the following sense. If
@@
with $\alpha_{u,i},\alpha_{v,i}\geqslant 0$, then the block $0^{\alpha_{u,i}}$ is translated to the block $0^{\alpha_{v,i+j}}$ (here we mean that the blocks could be empty).
In addition, if $u^{\prime}$ and $v^{\prime}$ are equal subwords of $u$ and $v$, respectively, and $n_{1,u^{\prime}}=n_{1,u}=n_{1,v}=n_{1,v^{\prime}}$, then we can overlay by $1$’s the word $u$ on $v$ such that $u^{\prime}$ and $v^{\prime}$ coincide. In other words, if
@@
We say that a subword $s$ of a cyclic word $w$ is unique if there is only one occurrence of $s$ in $w$, i.e., if we index 0’s and 1’s in $s$ and in $w$, then there is a unique injection from indices of letters in $s$ to indices of letters in $w$ giving an occurrence of $s$ in $w$.
@@
Let $w=01011=0_{1}1_{1}0_{2}1_{2}1_{3}$. Then the word $111$ is not unique, since we can take it in three ways $1_{1}1_{2}1_{3},1_{2}1_{3}1_{1},1_{3}1_{1}1_{2}$, and the word $0110$ is unique, since we can take in one way $0_{2}1_{2}1_{3}0_{1}$.
We remark that in a periodic cyclic word $w$ there are no unique subwords: indeed, we can shift the indices by the period. Similarly, a periodic subword of any cyclic word cannot be unique.
@@
Let $w$ be a cyclic word with indexed $1$’s: $w=1_{1}0^{\alpha_{1}}1_{2}0^{\alpha_{2}}1_{3}0^{\alpha_{3}}\cdots 1_{n_{1}}0^{\alpha_{n_{1}}}$ ($\alpha_{i}\geqslant 0$). A unit turn of $w$ is a word of the form $1_{1}0^{\alpha_{i}}1_{2}0^{\alpha_{i+1}}1_{3}0^{\alpha_{i+2}}\cdots 1_{n_{1}}0^{\alpha_{i-1}}$ for some $i$, where indices are taken modulo $n_{1}$ (from 1 to $n_{1}$).
Essentially, a unit turn of $w$ corresponds to overlay by 1’s from Definition 2.3 of a word with itself. Note that a unit turn of the word $w$ generates a unit turn of its subwords in the following sense. Let $w$ be a cyclic word with indexed $1$’s: $w=1_{1}0^{\alpha_{1}}1_{2}0^{\alpha_{2}}1_{3}0^{\alpha_{3}}\cdots 1_{n_{1}}0^{\alpha_{n_{1}}}$ ($\alpha_{i}\geqslant 0$). Let $s_{1}=1_{1}0^{\beta_{1}}1_{2}0^{\beta_{2}}1_{3}0^{\beta_{3}}\cdots 1_{n_{1}}0^{\beta_{n_{1}}}$, $\beta_{i}\leqslant\alpha_{i}$, and $s_{2}=1_{1}0^{\gamma_{1}}1_{2}0^{\gamma_{2}}1_{3}0^{\gamma_{3}}\cdots 1_{n_{1}}0^{\gamma_{n_{1}}}$, $\gamma_{i}\leqslant\alpha_{i}$ be two subwords of $w$ containing all 1’s. We say that $s_{1}$ and $s_{2}$ differ by a unit turn if there exists a unit turn of $w$ such that it generates a unit turn of $s_{1}$ which is equal to $s_{2}$, i.e., there exists $j$ such that $\beta_{i}=\gamma_{i+j}$ for each $i$ ($i+j$ is taken modulo $n_{1}$). We will say that the unit turn of $w$ translates the subword $s_{1}$ to the equal subword $s_{2}$.
In the proof of the main result, we will make use of the following notation:
@@
Let $u$ and $v$ be cyclic words of the same length, such that $n_{1,u}=n_{1,v}$ and $x_{u}=x_{v}=x$. We write that $u_{short}\cong v_{short}$ if $u_{short}=v_{short}$ and we can add one block $0^{x}$ to $u_{short}$ and one block $0^{x}$ to $v_{short}$ such that the obtained words are subwords of $u$ and $v$, respectively, and they are equal.
If there are no such blocks $0^{x}$, then $u_{short}\ncong v_{short}$, even if $u_{short}=v_{short}$.
@@
This notion will be used only for special words.
3 Several propositions
In the following proposition we show that for each $n\geqslant 7$ there are pairs of words for which sets of subwords of length at most $\frac{3}{4}n-\frac{3}{2}$ are equal.

### v2

@@ Abstract
The problem of reconstructing words from their subwords involves determining the minimum amount of information needed, such as multisets of scattered subwords of a specific length or the frequency of scattered subwords from a given set, in order to uniquely identify a word. In this paper we show that a cyclic word on a binary alphabet can be reconstructed by its scattered subwords of length $\frac{3}{4}n+4$, and for each $n$ one can find two cyclic words of length $n$ which have the same set of scattered subwords of length $\frac{3}{4}n-\frac{3}{2}$.
@@ 1 Introduction
The problem of reconstruction of words by their subwords is an important topic in combinatorics of words and related fields of mathematics and theoretical computer science. There are many variants of this problem: reconstruction of normal or cyclic words, from the sets of scattered or contiguous subwords, with or without multiplicities, restricting to a particular family of words and so on. It has been known for more than 40 years that one can reconstruct a word of length $n$ from the set of all its scattered subwords of length $\lfloor\frac{n-1}{2}\rfloor$, and this bound is optimal [8]. The same bound holds for contiguous subwords with multiplicities, while without multiplicities it is (n1)(n-1) [4].
The question of reconstruction of words of length $n$ by their scattered subwords of length $k$ with multiplicities turned out to be more complicated. One of the first results related to this question states that for words over a binary alphabet and for $k\geqslant\lfloor\frac{n}{2}\rfloor$ we can uniquely reconstruct the word, and for $k<\log_{2}(n)$ we cannot do it [5]. One of the best upper bounds of this problem with subword length $\lfloor\frac{16}{7}\sqrt{n}\rfloor$ follows from a result related to polynomials introduced in [1]. One of the best lower bounds is $k=\Omega(e^{\log^{\frac{1}{2}}(n)})$ [2].
Other results about words reconstruction include reconstruction from the number of occurrences of scattered factors of some special form. For example, a word $w\in\{a,b\}^{*}$ can be reconstructed from the number of occurrences of at most $\min(|w|_{a},|w|_{b})+1$ scattered factors of the form $a^{i}b$, where $|w|_{a}$ is the number of occurrences of the letter $a$ in $w$ [3]. A word can also be reconstructed from the number of occurrences as scattered subwords of $O(\ell^{2}\sqrt{n\log(n)})$ words we have chosen, where $\ell$ is the cardinality of alphabet [9] (this improves the results from [3]). Besides that, there are some results about reconstruction of words of special types. For example, reconstruction of DNA-words has been studied in [7], and reconstruction of words that are cyclic shifts of some fixed word has been explored in [6].
@@
Problem: Consider a cyclic binary word $w$ of length $n$ and the set of its scattered cyclic subwords of length at most $k$ (without multiplicities). For which $k$ can we always recover $w$?
@@
For any two distinct cyclic binary words $u$ and $v$ of length $n$ there exists a word $w$ of length at most $\frac{3}{4}n+4$ which is a subword of exactly one of the words $u$ and $v$.
We also show that for each $n\geqslant 7$ there are pairs of words for which sets of subwords of length at most $\frac{3}{4}n-\frac{3}{2}$ are equal (see Proposition 3.1). So, we found a lower bound $k\geqslant\frac{3}{4}n-\frac{3}{2}$ and an upper bound $k\leqslant\frac{3}{4}n+4$. In other words, we calculated $k$ modulo some constant which is at most $5$, and therefore our bound is almost sharp.
The paper is organized as follows. In Section 2, we introduce necessary definitions and notation used throughout the text. In Section 3, we prove some auxiliary propositions that are needed for the proof of the theorem. Besides that, in this section we provide a lower bound for $k$. In Section 4, we prove the main theorem. The section is divided into two subsections: in Subsection 4.1 we introduce the notation used throughout the section and the general structure of the proof, and Subsection 4.2 contains the proof of the theorem divided into five lemmas.
@@ 2 Definitions and notation
Two words $x$ and $y$ are said to be conjugate if there exist words $u$ and $v$ such that $x=uv$ and $y=vu$. We define a cyclic word as an equivalence class of the conjugacy relation on $\Sigma^{*}$. Thus, if $w\in\Sigma^{*}$, then the cyclic word represented by $w$ is the set $\{vu\in\Sigma^{*}\mid u,v\in\Sigma^{*},uv=w\}$. For the rest of this paper, when referring to a cyclic word, we write a representative of this class, slightly abusing the notion to avoid cluttering the text. For a finite or a cyclic word its length is the number of letters in it.
For a cyclic word $w$ with a representative $w_{1}\cdots w_{n}$, a subword of $w$ is a cyclic word with a representative of the form $w_{i_{1}}\cdots w_{i_{k}}$, where $1\leqslant i_{1}<\ldots<i_{k}\leqslant n$. A factor of some conjugate of $w_{1}\cdots w_{n}$ of the form $0^{+}$ (resp., $1^{+}$) continued to the left and to the right with 1 (resp., 0) is called a block of 0’s (resp., 1’s).
For two cyclic words $u$ and $v$, we say that a cyclic word $w$ is a distinguishing subword if it is a subword of only one of the words $u$ and $v$. Using this notion, Theorem 1.1 can be reformulated as follows: any two distinct cyclic binary words of length $n$ have a distinguishing subword of length at most $\frac{3}{4}n+4$.
We say that a cyclic word $w$ is periodic if $w=(1^{\alpha_{1}}0^{\beta_{1}}1^{\alpha_{2}}0^{\beta_{2}}\cdots 1^{\alpha_{s}}0^{\beta_{s}})^{r}$ for some $\alpha_{i}\geqslant 1$, $\beta_{i}\geqslant 1$, $s\geq 1$ and $r\geqslant 2$.
Let $w$ be a cyclic word of length $n$. Throughout the paper, we make use of the following notation:
  • $n_{0,w}$ and $n_{1,w}$ are the numbers of 0’s and 1’s in $w$.
  • $2l_{w}$ is the total number of blocks in $w$, $l_{w}$ blocks of 0’s and $l_{w}$ blocks of 1’s (here we consider a representative of $w$ in which the first and the last letters are distinct).
  • $x_{w}$ is the length of the longest block of 0’s in $w$.
@@
  • $y_{w}$ is the length of the longest block in $w$ which is smaller than $x_{w}$ (we might have $y_{w}=0$).
  • $n_{\operatorname{long},w}$ is the number of 0’s in big blocks $0^{x_{w}}$ and $n_{\operatorname{short},w}=n_{0,w}-n_{\operatorname{long},w}$ is the number of 0’s in small blocks.
  • $w_{\operatorname{long}}$ is the subword of $w$ which contains all 1’s of $w$ and all 0’s from big blocks of 0’s (all blocks $0^{x_{w}}$).
  • $w_{\operatorname{short}}$ is the subword of $w$ which contains all 1’s of $w$ and all 0’s from small blocks of 0’s (all blocks of 0’s except for the blocks $0^{x_{w}}$).
  • $w_{\operatorname{zeros}}$ is the subword of $w$ which contains all 1’s of $w$ and one 0 from each block of 0’s.
For the rest of this paper, we omit the subscript $w$ when no confusion arises. The notation $n_{0},n_{1}$ is justified by the following. In the beginning of the next section we prove Proposition 3.3 stating that cyclic words $u$ and $v$ with the same set of subwords of length $\frac{3}{4}n+4$ have the same numbers of 0’s, 1’s and also the same number of blocks. So, for the rest of the paper $n_{0},n_{1}$ and $l$ are fixed.
@@
Let $w=10^{3}101^{2}0^{3}1^{2}0^{2}$. Then $n_{1}=6,n_{0}=9,x=3,n_{\operatorname{long}}=6,n_{\operatorname{short}}=3,w_{\operatorname{long}}=1^{3}0^{3}1^{3}0^{3},w_{\operatorname{zeros}}=10101^{2}01^{2}0$.
For the proof of the main result, we need to treat words of certain specific forms separately. We hence introduce the following definition:
@@
Let $w$ be a word with $l_{w}\geqslant 2$. We say that $w$ is special if it is of one of the following three types:
@@
    $w=(0^{t}1^{m})^{i}0^{t}1^{2m}(0^{t}1^{m})^{l_{w}-i-2}0^{t}1^{2m}$
    for some positive integers $t$, $m$ and $i$.
So, a word of the first type has $l_{w}$ blocks $1^{m}$, a word of the second type has $l_{w}-1$ blocks $1^{m}$ and one block $1^{2m}$, a word of the third type has $l_{w}-2$ blocks $1^{m}$ and two blocks $1^{2m}$, and the number of blocks of 0’s is at least two in each case.
For a special word $w$, we let $dist(w)$ denote the length of the shortest block of 1’s, i.e., the minimum distance between blocks of 0’s. Note that the value of $dist(w)$ is equal to the number $m$ from the definition of a special word.
When working with cyclic words, sometimes we need to index their letters, either just one letter or both letters. For example, for a cyclic word generated by $00101$ we could either index 1’s as $001_{1}01_{2}$, or as $001_{2}01_{1}$. We remark that it is not exactly the same as choosing a representative from the conjugacy class (the difference comes up in periodic cyclic words).
@@
Let $u$ and $v$ be cyclic words with indexed 1’s such that $n_{0,u}=n_{0,v}$ and $n_{1,u}=n_{1,v}$. We then define a 1-overlay of the word $u$ on the word $v$ as a bijection between indexed 1’s in $u$ and in $v$ which, for some integer $i$, translates each $1_{j}$ in $u$ to $1_{j+i}$ in $v$ (indices are taken modulo $n_{1}$).
An example of a $1$-overlay is provided on Fig. 1.
[…]
Figure 1: Example for Definition 2.3: $u=1_{1}0^{3}1_{2}1_{3}0^{2}1_{4}0^{2}$ (left), $v=1_{1}01_{2}0^{3}1_{3}1_{4}0^{3}$ (center), 1-overlay of $u$ on $v$ for $i=2$ (right).
In other words, a 1-overlay can be seen as an order-preserving bijection between 1’s in the words $u$ and $v$, or simply as a shift of indices of 1’s in $v$ relative to in $u$. This bijection between 1’s induces a bijection between blocks of 0’s in the following sense. If
@@
with $\alpha_{u,i},\alpha_{v,i}\geqslant 0$, then $0^{\alpha_{u,i}}$ is translated to $0^{\alpha_{v,i+j}}$. We remark that $\alpha_{u,i}$ and $\alpha_{v,i+j}$ can be equal to 0. If $\alpha=0$, we say that $0^{\alpha}$ is empty place or $\varnothing$ .
In addition, if $u^{\prime}$ and $v^{\prime}$ are equal subwords of $u$ and $v$, respectively, and $n_{1,u^{\prime}}=n_{1,u}=n_{1,v}=n_{1,v^{\prime}}$, then we can consider a 1-overlay of $u$ on $v$ such that $u^{\prime}$ and $v^{\prime}$ coincide. In other words, if
@@
We say that a subword $s$ of a cyclic word $w$ is unioccurrent if there is only one occurrence of $s$ in $w$, i.e., if we index 0’s and 1’s in $s$ and in $w$, then there is a unique injection from indices of letters in $s$ to indices of letters in $w$ giving an occurrence of $s$ in $w$.
@@
Let $w=01011=0_{1}1_{1}0_{2}1_{2}1_{3}$. Then the word $111$ is not unioccurrent, since we can choose it in three ways $1_{1}1_{2}1_{3},1_{2}1_{3}1_{1},1_{3}1_{1}1_{2}$, and the word $0110$ is unioccurrent, since we can take in one way $0_{2}1_{2}1_{3}0_{1}$.
We remark that in a periodic cyclic word $w$ there are no unioccurrent subwords: indeed, we can shift the indices by the period. Similarly, a periodic subword of any cyclic word cannot be unioccurrent.
@@
Let $w$ be a cyclic word with indexed 1’s: $w=1_{1}0^{\alpha_{1}}1_{2}0^{\alpha_{2}}1_{3}0^{\alpha_{3}}\cdots 1_{n_{1}}0^{\alpha_{n_{1}}}$ ($\alpha_{i}\geqslant 0$). A turn of $w$ is a word of the form $1_{1}0^{\alpha_{i}}1_{2}0^{\alpha_{i+1}}1_{3}0^{\alpha_{i+2}}\cdots 1_{n_{1}}0^{\alpha_{i-1}}$ for some $i$, where indices are taken modulo $n_{1}$ (from 1 to $n_{1}$).
An example illustrating the above definition is provided on Fig. 2.
$_{2}11$$_{1}$$1_{3}$$1_{4}$$^{3}00$$0^{2}$$_{2}11$$_{1}$$1_{3}$$1_{4}$$0^{2}0$$^{3}0$$_{2}11$$_{1}$$1_{3}$$1_{4}$$^{2}00$$0^{3}$$_{2}11$$_{1}$$1_{3}$$1_{4}$$0^{3}0$$^{2}0$
Figure 2: Example for Definition 2.6: four turns of the word $1_{1}0^{3}1_{2}01_{3}0^{2}1_{4}0$.
Essentially, a turn of $w$ corresponds to a 1-overlay from Definition 2.3 of a word with itself. Note that a turn of the word $w$ generates a turn of its subwords in the following sense. Let $w$ be a cyclic word with indexed 1’s: $w=1_{1}0^{\alpha_{1}}1_{2}0^{\alpha_{2}}1_{3}0^{\alpha_{3}}\cdots 1_{n_{1}}0^{\alpha_{n_{1}}}$ ($\alpha_{i}\geqslant 0$). Let $s_{1}=1_{1}0^{\beta_{1}}1_{2}0^{\beta_{2}}1_{3}0^{\beta_{3}}\cdots 1_{n_{1}}0^{\beta_{n_{1}}}$, $\beta_{i}\leqslant\alpha_{i}$, and $s_{2}=1_{1}0^{\gamma_{1}}1_{2}0^{\gamma_{2}}1_{3}0^{\gamma_{3}}\cdots 1_{n_{1}}0^{\gamma_{n_{1}}}$, $\gamma_{i}\leqslant\alpha_{i}$ be two subwords of $w$ containing all 1’s. We say that $s_{1}$ and $s_{2}$ differ by a turn if there exists a turn of $w$ such that it generates a turn of $s_{1}$ which is equal to $s_{2}$, i.e., there exists $j$ such that $\beta_{i}=\gamma_{i+j}$ for each $i$ ($i+j$ is taken modulo $n_{1}$). We further say that the turn of $w$ translates the subword $s_{1}$ to the equal subword $s_{2}$. We would like to emphasize that for the notion of a turn, letters are not treated symmetrically.
In the proof of the main result, we make use of the following notation:
@@
Let $u$ and $v$ be cyclic words of equal length, such that $n_{1,u}=n_{1,v}$ and $x_{u}=x_{v}=x$. Let $u_{1}$ and $v_{1}$ be subwords of $u_{short}$ and $v_{short}$ correspondingly. We write $u_{1}\cong v_{1}$ if $u_{1}=v_{1}$ and we can add one block $0^{x}$ to $u_{1}$ and one block $0^{x}$ to $v_{1}$ such that the obtained words are subwords of $u$ and $v$, respectively, and they are equal.
Note that if there are no such blocks $0^{x}$, then $u_{\operatorname{short}}\ncong v_{\operatorname{short}}$, even if $u_{\operatorname{short}}=v_{\operatorname{short}}$.
Example 2.8.
Let $x>1$ and $u=0^{x}10110^{x}1011111$, $v=0^{x}101110110^{x}111$ and $w=0^{x}11010^{x}1101111$. Then $u_{\operatorname{short}}=v_{\operatorname{short}}=w_{\operatorname{short}}=10111011111$ and $u_{\operatorname{short}}\cong v_{\operatorname{short}}$, since we can add a block $0^{x}$ to $u_{\operatorname{short}}$ and add a block $0^{x}$ to $v_{\operatorname{short}}$ and get the word $0^{x}10111011111$. However, we have $u_{\operatorname{short}}\ncong w_{\operatorname{short}}$ and $v_{\operatorname{short}}\ncong w_{\operatorname{short}}$.
@@
This notion is used only for special words.
3 Preliminary observations and auxiliary statements
In the following proposition, we show that for each $n\geqslant 7$ there are pairs of words for which the sets of their subwords of length at most $\frac{3}{4}n-\frac{3}{2}$ are equal.